1. (a) Use the Mean Value Theorem to establish the following inequality

e a (x − a) < ex − e a < ex (x − a), if a < x.




1
Expert's answer
2021-05-24T16:05:20-0400

1.(a)


(x-a) < ex - e a < e (x-a) if a<x

Let f(x) = ex

f'(x) = ex



From Mean Value Theorem,


"\dfrac{f(b) - f(a)}{ (b-a)} = f'(c)" ∀ c ∈ (a,b)


let a < x < b


so


"\dfrac{e^x - e^ a }{(x-a)} = e ^c" .............................(1)


a < c < x

< e < ex

(x-a) < ex - e a < e (x-a) ..................from equation (1)